Need help with this word problem.
4) CLASSIC PROBLEM - A traveling salesman (selling shoes) stops at a farm in the Midwest. Before he could knock on the door, he noticed an old truck on fire. He rushed over and pulled a young lady out of the flaming truck. Farmer Crane came out and gratefully thanked the traveling salesman for saving his daughter’s life. Mr. Crane insisted on giving the man an award for his heroism.
So, the salesman said, “If you insist, I do not want much. Get your checkerboard and place one grain of wheat on the first square. Then place two grains of wheat on the next square. Then place four grains on the third square. Continue this until all 64 squares are covered with grains of wheat.” As he had just harvested his wheat, Mr. Crane did not consider this much of an award, but he soon realized he made a miscalculation on the amount of wheat involved.
a) How much wheat would Mr. Brown have to put on the 24nd square?
Answer:
Show work in this space.

Answer by fractalier(2101)

Yes, a classic problem, this time presented as grains of wheat. I first heard it as grains of rice given by the Chinese emperor to a great warrior...
This is a "powers of two" problem...
1st square = 2^0 grains
2nd square = 2^1 grains
3rd square = 2^2 grains
4th square = 2^3 grains
.
.
.
24th square = 2^23 grains ( a very large number )